What will be the energy of the 5th orbit of hydrogen atom if the energy of the 1st orbit of this atom is 13.6 eV?


1
Expert's answer
2021-06-21T15:05:15-0400

Gives

energy of "1^{st}" orbital H-atom

"E=13.6ev"

n=5

Now

We know that



"E_n=-\frac{-13.6}{n^2}=\frac{-13.6}{5^2}=-0.544eV"

"E_5=-0.544eV"